The height h in feet of an object after t seconds is given by the function
h = -16t^2 + 40t + 10.
How high was the object when it was launched?

Answer by ikleyn(52781)

.

The initial height is the value of the function h(t) at t = 0


    h(0) = -16*0^2 + 40*0 + 10 = 0 + 0 + 10 = 10.


ANSWER.  Initial height is 10 feet.

Solved, answered and explained.